# Fine-Tuning and Prompt Engineering for Large Language Models-based Code Review Automation

### v3

@@ 3.3 Model Fine-Tuning
Figure 3: Prompt templates for zero-shot learning and few-shot learning (lang refers to a programming language). The text in blue is omitted when reviewers’ comments are not used in experiments.
@@ 4 Result
Approach. To address this RQ, we leverage fine-tuning and inference techniques (i.e., zero-shot learning and few-shot learning) on GPT-3.5 and Magicoder (The details of GPT-3.5 and Magicoder are presented in Table 5) . Then, we measure EM of the results obtained from GPT-3.5, Magicoder and Guo et al.’s approach [14].
@@
Result. The fine-tuning of GPT 3.5 with zero-shot learning helps GPT-3.5 to achieve 73.17% -74.23% higher EM than the Guo et al. [14]’s approach. Table 4 shows the results of EM achieved by GPT-3.5, Magicoder and Guo et al.’s approach [14]. The table shows that when GPT-3.5 and Magicoder are fine-tuned, such models achieves 73.17% -74.23% and 26.00% - 28.53% higher EM than the Guo et al.’s approach [14], respectively.
@@
When a persona is included in input prompts, GPT-3.5 achieves 1.02% - 54.17% lower EM than when the persona is not included in input prompts. Table 4 also shows that in terms of EM, GPT-3.5 that zero-shot and few-shot learning with a persona are used achieves 3.67% - 54.17% and 1.02% - 30.77% lower than GPT-3.5 that zero-shot and few-shot learning without a persona are used, respectively. In terms of CodeBLEU, GPT-3.5 that zero-shot and few-shot learning with a persona are used achieves 1.33% - 19.13% and 0.15% - 0.90% lower than GPT-3.5 that zero-shot and few-shot learning without a persona are used, respectively.
The results indicate that GPT-3.5 that zero-shot learning and few-shot learning with a persona are used achieves lower EM and CodeBLEU. To illustrate the impact of the persona, we present an example of the revised code that GPT-3.5 generates when zero-shot learning with and without a persona are used, and an example of the revised code that GPT-3.5 generates when few-shot learning with and without a persona are used in Figure 4.
5 Discussion
In this section, we discuss the implications of our findings, the additional results of GPT-3.5., and the cost and benefits of using GPT-3.5.
@@ 5.1 Implications of Our Findings
GPT-3.5 does not require a lot of training data for model fine-tuning to adapt to the code review automation task since the results of RQ1 show that GPT-3.5 that is fine-tuned on a subset of a training set outperforms the studied code review automation approaches [4, 5, 6]. The results imply that GPT-3.5 can adapt to the code review automation task by learning from a small set of training examples (approximately 20k training examples in this study), unlike the studied code review automation approaches that require the whole training set to adapt to the code review automation task.
@@
Figure 6 presents the characteristic of the code changes (i.e., Fixing Bug, Refactoring and Other) of the revised code that GPT-3.5 correctly generates. According to the figure, for the Tufanodata (with comment), CodeReviewerdata and D-ACTdata dataset, the fine-tuning of GPT 3.5 with zero-shot learning helps GPT-3.5 achieves the highest EM for the code changes of Refactoring and Other. In contrast, for the Tufanodata (without comment) dataset, non fine-tuned GPT-3.5 with few-shot learning achieves the highest EM for the code changes of all categories. According to Figure 6(a) and Figure 6(b), we also find that fine-tuned GPT-3.5 and non fine-tuned GPT-3.5 that zero-shot and few-shot learning are used achieve the highest EM for the code changes of type other across all studied datasets. The reason for this result is that we do not specify the characteristics of the revised code (i.e., fixing bug and refactoring) in prompts. Thus, such models possibly generate revised code that is not specific to fixing bug or refactoring. Therefore, the majority of the code changes of the generated revised code are categorized as other.
@@ 6.1 Threats to Construct Validity
Threats to construct validity relate to the example selection techniques that we use to select examples for fine-tuning GPT-3.5 and doing few-shot learning; and the design choices of our prompt. We explain each threat below.
In our study, we randomly select 6% examples from a training set to fine-tune GPT-3.5. We do not use the whole training set since it is prohibitively expensive (i.e., it could cost more than hundreds of dollars to fine-tune a single model). We use only 6% examples of a training set to fine-tune GPT-3.5 since we find that the fine-tuned GPT-3.5 outperforms the existing code review automation approaches. However, we do not conduct experiments by using less than 6% examples from a training set to fine-tune GPT-3.5 since the main goal of our experiment is not to find the smallest set of examples from a training set to fine-tune GPT-3.5 so that fine-tuned GPT-3.5 outperforms the existing code review automation approaches. In contrast, the main goal of our experiment is to demonstrate that using a subset of examples from a training set to fine-tune GPT-3.5 could help fine-tuned GPT-3.5 outperform the existing code review automation approaches. Thus, the relationship between the number of selected examples from a training set and the performance of fine-tuned GPT-3.5 can be explored in future work.
In addition, we only use the BM25 [41] technique to select three demonstration examples for few-shot learning. However, using more demonstration examples or different techniques to select demonstration examples may lead to results that differ from the reported results. Thus, more demonstration examples and other techniques for selecting demonstration examples can be explored in future work.
Another threat is the design choice of our prompts may be sub-optimal. To mitigate this threat, we follow the guidelines of prompt design from OpenAI. However, future work could experiment with other prompt designs to find the most suitable prompts for code review automation.
Finally, we use only two prompt templates (one for zero-shot learning and another one for few-shot learning) when conducting experiments. We do not explore multiple prompts since doing so can be expensive (i.e., it can cost nearly a thousand dollars to conduct experiments with all datasets and settings for one prompt template). In addition, since the code review automation task is related to revising the patches written by software developers, we use the persona (i.e., an expert software developer) to ensure that the revised code generated by GPT-3.5 looks like the source code written by a software developer. However, there are other similar personas (e.g., a senior software engineer, or a front-end software developer) that we do not explore. Thus, future work can explore other design choices of prompt and persona to find the optimal prompt and persona for code review automation tasks.

### v4

@@ 3.3 Model Fine-Tuning
Figure 3: Prompt templates for zero-shot learning and few-shot learning that contain simple instructions (lang refers to a programming language). The text in blue is omitted when reviewers’ comments are not used in experiments.
@@ 4 Result
Approach. To address this RQ, we leverage fine-tuning and inference techniques (i.e., zero-shot learning, few-shot learning, and persona) on GPT-3.5 and Magicoder (The details of GPT-3.5 and Magicoder are presented in Table 5) . Then, we measure EM of the results obtained from GPT-3.5, Magicoder and Guo et al.’s approach [14].
@@
Result. The fine-tuning of GPT 3.5 with zero-shot learning helps GPT-3.5 to achieve 73.17% -74.23% higher EM than the Guo et al. [14]’s approach. Table 4 shows the results of EM achieved by GPT-3.5, Magicoder and Guo et al.’s approach [14]. The table shows that when GPT-3.5 and Magicoder are fine-tuned, such models achieve 73.17% -74.23% and 26.00% - 28.53% higher EM than the Guo et al.’s approach [14], respectively.
@@
When a persona is included in input prompts, GPT-3.5 achieves 1.02% - 54.17% lower EM than when the persona is not included in input prompts. Table 4 also shows that when a persona is included in prompts, GPT-3.5 with zero-shot and few-shot learning achieves 3.67% - 54.17% and 1.02% - 30.77% lower EM compared to when a persona is not included in prompts, respectively. Similarly, when a persona is included in prompts, GPT-3.5 with zero-shot and few-shot learning achieves 1.33% - 19.13% and 0.15% - 0.90% lower CodeBLEU compared to when a persona is not included in prompts, respectively.
The results indicate that when a persona is included in prompts, GPT-3.5 with zero-shot and few-shot learning achieves lower EM and CodeBLEU. To illustrate the impact of the persona, we present an example of the revised code that GPT-3.5 generates when zero-shot learning with and without a persona is used, and an example of the revised code that GPT-3.5 generates when few-shot learning with and without a persona is used in Figure 4.
5 Discussion
In this section, we discuss the implications of our findings, the additional results of GPT-3.5, and the cost and benefits of using GPT-3.5.
@@ 5.1 Implications of Our Findings
GPT-3.5 does not require a lot of training data for model fine-tuning to adapt to the code review automation task since Table 4 shows that GPT-3.5 that is fine-tuned on a subset of a training set outperforms the studied code review automation approaches [4, 5, 6]. The results imply that GPT-3.5 can adapt to the code review automation task by learning from a small set of training examples (approximately 20k training examples in this study), unlike the studied code review automation approaches that require the whole training set to adapt to the code review automation task.
@@
Figure 6 presents the characteristics of the code changes (i.e., Fixing Bug, Refactoring and Other) of the revised code that GPT-3.5 correctly generates. According to the figure, for the Tufanodata (with comment), CodeReviewerdata and D-ACTdata dataset, the fine-tuning of GPT 3.5 with zero-shot learning helps GPT-3.5 achieves the highest EM for the code changes of Refactoring and Other. In contrast, for the Tufanodata (without comment) dataset, non fine-tuned GPT-3.5 with few-shot learning achieves the highest EM for the code changes of all categories. According to Figure 6(a) and Figure 6(b), we also find that fine-tuned GPT-3.5 and non fine-tuned GPT-3.5 that zero-shot and few-shot learning are used achieve the highest EM for the code changes of type other across all studied datasets. The reason for this result is that we do not specify the characteristics of the revised code (i.e., fixing bug and refactoring) in prompts. Thus, such models possibly generate revised code that is not specific to fixing bug or refactoring. Therefore, the majority of the code changes of the generated revised code are categorized as other.
Table 6: The evaluation results of GPT-3.5 when being fine-tuned with different sizes of training sets.
CodeReviewer Tufano (with comment) Tufano (without comment) Android Google Ovirt
Size of Training Set EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU
6% 37.93% 49.00% 22.16% 82.99% 6.02% 79.81% 2.34% 74.15% 6.71% 81.08% 3.05% 74.67%
10% 37.72% 48.83% 22.31% 83.43% 5.37% 80.68% 2.51% 75.10% 5.98% 80.65% 2.71% 75.13%
20% 38.80% 49.33% 22.84% 83.44% 5.65% 80.42% 2.34% 76.04% 7.52% 81.40% 2.83% 75.46%
5.3 The Impact of the Size of Training Dataset on Fine-Tuned GPT-3.5
The results of RQ2 show that model fine-tuning can help increase the performance of GPT-3.5. However, little is known whether fine-tuned GPT-3.5 can achieve higher performance when being fine-tuned with larger training sets. Therefore, we conduct experiments by using 10% and 20% of training sets to fine-tune GPT-3.5 (we do not use persona in these experiments).
Table 6 shows the results of EM and CodeBLEU that fine-tuned GPT-3.5 achieves across different sizes of training sets. The table shows that GPT-3.5 that is fine-tuned with 20% of a training set achieves 2.29% - 12.07% higher EM and 0.54% - 2.55% higher CodeBLEU than GPT-3.5 that is fine-tuned with 6% of a training set. In addition, GPT-3.5 that is fine-tuned with 20% of a training set achieves 2.38% - 25.75% higher EM and 0.01% - 1.25% higher CodeBLEU than GPT-3.5 that is fine-tuned with 10% of a training set, respectively. The results indicate that fine-tuned GPT-3.5 achieves higher performance when being fine-tuned with a larger training set.
(a) A prompt template for zero-shot learning.
(b) A prompt template for few-shot learning.
Figure 7: Prompt templates for zero-shot learning and few-shot learning that the instructions are broken into smaller tasks (lang refers to a programming language). The text in blue is omitted when reviewers’ comments are not used in experiments.
(a) A prompt template for zero-shot learning.
(b) A prompt template for few-shot learning.
Figure 8: Prompt templates for zero-shot learning and few-shot learning that more details are added to the instructions (lang refers to a programming language). The text in blue is omitted when reviewers’ comments are not used in experiments.
Table 7: The evaluation results of GPT-3.5 for different prompt templates. P1 refers to the prompt templates with simple instructions (Figure 3). P2 refers to the prompt templates with instructions being broken down into smaller steps.(Figure 7). P3 refers to the prompt templates with detailed instructions (Figure 8).
CodeReviewer Tufano (with comment) Tufano (without comment) Android Google Ovirt
Prompt Design Prompting EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU EM CodeBLEU
P1 Zero-shot 17.72% 44.17% 13.52% 78.36% 2.62% 74.92% 0.49% 61.85% 0.16% 61.04% 0.48% 56.55%
P2 14.47% 43.52% 11.24% 79.05% 2.25% 76.54% 0.54% 66.10% 0.16% 67.07% 0.33% 60.76%
P3 11.94% 41.18% 9.86% 76.18% 1.26% 72.31% 0.05% 53.03% 0.08% 46.22% 0.26% 41.20%
P1 Few-shot 26.55% 47.50% 19.79% 81.47% 8.96% 79.21% 2.34% 75.33% 2.89% 81.40% 1.64% 73.83%
P2 25.25% 48.45% 15.82% 80.16% 6.84% 76.50% 0.60% 75.94% 3.56% 81.40% 1.67% 74.18%
P3 25.14% 48.60% 15.07% 79.83% 5.81% 75.76% 0.38% 74.95% 2.91% 81.18% 1.49% 73.31%
5.4 The Impact of Prompt Design on GPT-3.5
In RQ3, we use the prompt templates in Figure 3 that contain simple instructions to conduct experiments. However, prior work [44, 45] found that the design of prompts has an impact on the performance of LLMs. Thus, we further investigate the impact of prompt design on GPT-3.5 for code review automation. To do so, we conduct experiments by using the following two new prompt designs (we do not include a persona in these prompt designs since the results of RQ3 show that GPT-3.5 without a persona outperforms GPT-3.5 with a persona). First, we use the prompt design that a single instruction is broken into smaller steps [^7] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/advanced-prompt-engineering?pivots=programming-language-chat-completions#break-the-task-down, as depicted in Figure 7. Second, we use the prompt design that contains more detailed instructions [^8] https://www.promptingguide.ai/introduction/tips, as depicted in Figure 8.
Table 7 shows the results of EM and CodeBLEU that GPT-3.5 achieves across different prompt designs. The table shows that for zero-shot learning, GPT-3.5 that is prompted by the prompt with a simple instruction achieves 16.44% - 45.45% higher EM than GPT-3.5 that is prompted by the prompt with an instruction being broken down into smaller steps. In addition, GPT-3.5 that is prompted by the prompt with a simple instruction achieves 37.12% - 880.00% higher EM than GPT-3.5 that is prompted by the prompt with a detailed instruction.
The table also shows that for few-shot learning, GPT-3.5 that is prompted by the prompt with a simple instruction achieves 5.15% - 290.00% higher EM than GPT-3.5 that is prompted by the prompt with an instruction being broken down into smaller steps. Furthermore, GPT-3.5 that is prompted by the prompt with a simple instruction achieves 5.61% - 515.79% higher EM than GPT-3.5 that is prompted by the prompt with detailed instruction.
The results indicate that GPT-3.5 that is prompted by the prompts with a simple instruction achieves the highest EM when compared to other prompt designs. Thus, the results imply that the prompt with a simple instruction is the most suitable for GPT-3.5 for code review automation.
@@ 6.1 Threats to Construct Validity
Threats to construct validity relate to the example selection techniques that we use to select examples for few-shot learning, and the design choices of the persona. We explain each threat below.
In this study, we only use the BM25 [41] technique to select three demonstration examples for few-shot learning. However, using more demonstration examples or different techniques to select demonstration examples may lead to results that differ from the reported results. Thus, more demonstration examples and other techniques for selecting demonstration examples can be explored in future work.
Since the code review automation task is related to revising the patches written by software developers, we use the persona (i.e., an expert software developer) to ensure that the revised code generated by GPT-3.5 looks like the source code written by a software developer. However, there are other similar personas (e.g., a senior software engineer, or a front-end software developer) that we do not explore. Thus, future work can explore other design choices of prompt and persona to find the optimal prompt and persona for code review automation tasks.